OpenStudy (anonymous):

The leanth of a rectangle is 5 centimeters less than twice its width. The perimter of the rectangle is 26cm.What are the dimensions?

OpenStudy (akashdeepdeb):

Length = 2*width - 5 l = 2w - 5 And we given is: perimeter of rectangle = 2*(length + width) = 26 Thus, 2(l+w) = 26 l+w = 13 2w-5 + w = 13 [ because l = 2w - 5 ] 3w = 18 thus, w = 6 Now get the value of l = 2*w - 5 from w's value. Understood? :)

OpenStudy (anonymous):

so the length would be 11 right ??

OpenStudy (anonymous):

I meant to say 7

OpenStudy (akashdeepdeb):

correct. :D